The height of a person is a random variable with variance < 5 inches. According to Mr. Chebyshev, how many people do we need to sample to ensure that the sample mean is at most 1 inch away from the distribution mean with probability 2 95%?
Answer #1

P(|X - mean| < k ) > 1 - \sigma^2/k^2

here replace X with Xbar

P(|Xbar - mean| < k)>= 1 - \sigma^2/(n*k^2 )

put k = 1

\sigma^2= 5

1- \sigma^2/(n k^2) = 0.95

1-5/n = 0.95

5/n = 0.05

n = 100